# On Selection of Cross-Section Averages in Non-stationary Environments

### v1

@@ 2 Econometric Setup
The discussion above seems to suggest that incorporating the degree of non-stationarity of $\mathbf{F}$ into the objective function should solve the issue and allow keeping $p_{N,T}$ the same. However, this is not the case. Let us assume that $\tau$ is known and normalize the objective function by $T^{1+\tau}$, instead, as suggested by Assumption 1 (a). In the Supplement, we show that
@@
$\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}=\ln\mathrm{det}\left[\mathbf{I}_{k}+\underbrace{\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x}}}}-\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}})\mathbf{X}_{i}}_{O_{p}(1)}\underbrace{\left(\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}}\mathbf{X}_{i}\right)^{-1}}_{O_{p}(T^{\tau})}\right],$ (2.6)
which means that, again, the determinant does not converge to a strictly positive constant. The reason behind this is that the rank condition is satisfied in the second component of the product, and by following steps of Karabiyik et al., (2017) and Westerlund, (2018), we obtain
$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}}\mathbf{X}_{i}=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}+o_{p}(1)=O_{p}(T^{-\tau}),$ (2.7)
since the idiosyncratics are stationary. In effect, the new normalization induces imbalance elsewhere. Therefore, IC fundamentally depend not only on stationarity of factors, but also on stationarity of idiosyncratics. We confirm this in the additional Monte Carlo simulations presented in the Supplement, where the adjustment of the objective function becomes more harmful with $\tau$.
@@ 3 Simulation Study
The data generating process of the simulation follows Margaritella and Westerlund, (2023):
@@
where $\mathbf{X}_{i}=\left[\mathbf{x}_{i,1},...,\mathbf{x}_{i,T}\right]^{\prime}$ is a $T\times k$ matrix of observable regressors, $\boldsymbol{\beta}=\iota^{\prime}0.5$ a $k\times 1$ vector of unobserved parameters set to 0.5. $\mathbf{F}$ is a $T\times m$ matrix of $m$ unobserved common factors. For simplicity, we assume that $\mathbf{X}_{i}$ and $\mathbf{u}_{i}$ load on $m$ factors. The factor loadings are generated as $\boldsymbol{\Gamma}_{i}=[\mathbf{I}_{m},\mathbf{0}_{m\times(k-m)}]\psi_{i}$, that is, a matrix $m\times k$, with $\psi_{i}\sim N(1,1)$ and the elements of the vector $1\times m$ $\boldsymbol{\gamma}_{i}$ are drawn from $N(1,1)$.
In the case of time- and space-independent errors, $\mathbf{e}_{i,t}$ is drawn from $N(\mathbf{0}_{k+1},\mathbf{I}_{k+1})$. In the case of errors correlated weakly across time and cross-sections we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
$\mathbf{e}_{i,t}=(\epsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}=\mathbf{P}_{\rho}\mathbf{e}_{i,t-1}+\mathbf{z}_{\mathbf{w},i,t},\quad\mathbf{P}_{\rho}=\begin{bmatrix}\rho&\mathbf{0}_{1\times k}\\ \mathbf{0}_{k\times 1}&\rho_{v}\mathbf{I}_{k}\end{bmatrix},\quad\mathbf{z}_{\mathbf{w},i,t}=\begin{bmatrix}\sqrt{\frac{m\left(1-\rho^{2}\right)}{1+2J\kappa^{2}}}\boldsymbol{\epsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa\mathbf{w}_{i}^{\prime}\right)\\ \sqrt{\frac{\left(1-\rho_{v}^{2}\right)}{1+2J_{v}\kappa_{v}^{2}}}\boldsymbol{\Upsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa_{v}\boldsymbol{\tilde{\mathbf{w}}}_{i}^{\prime}\right)\end{bmatrix}$ (3.2)
where $\boldsymbol{\epsilon}_{t}$ is a $N\times 1$ stack of $\epsilon_{i,t}\sim N(0,1)$ over $i$, and $\mathbf{w}_{i}$ is the $i$-th row ($1\times N$) of $\mathbf{W}$, a weight matrix with the $J$-th off diagonal elements equal to zero. $\rho$ is set to 0.5, $\kappa=0.2$ and $J=5$. The idiosyncratics in $\mathbf{X}_{i}$, $\mathbf{V}_{i}$, are also drawn as potentially weakly correlated over time and cross-sections. Here, $\mathbf{\Upsilon}_{t}$ is a $N\times k$ matrix, which stacks $\boldsymbol{\upsilon}_{i,t}\sim N(\mathbf{0}_{k},\mathbf{I}_{k})$ over $i$, and it is independent of $\boldsymbol{\epsilon}_{t}$. The vector $\boldsymbol{\iota}_{N,i}\in\mathbb{R}^{N\times 1}$ contains 1 in the $i$-th coordinate, and zeros elsewhere. $\rho_{v}=0$ for the time IID case and $\rho_{v}=0.5$ for weakly time-correlated idiosyncratics. In both cases, we allow for weak correlation across cross sections, $\kappa_{v}=0.2$, $J_{v}=5$ and $\tilde{\mathbf{w}}_{i}$ is defined as $\mathbf{w}_{i}$ above. Therefore, we also test correlated idiosyncratics over time, whereas our theory prescribes uncorrelated innovations for $\tau\in(0,1)$. The factors are drawn as:
$\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{Q}T^{-\tau},\quad\mathbf{Q}=diag(q_{1},...,q_{m}),q_{j}\sim U(0,2)$ (3.3)
$\mathbf{u}_{t}\sim N\left(\mathbf{0}_{m},\sqrt{\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}}\right).$ (3.4)
@@
Table 3 below presents the average of the number of CAs selected for the case of stationary factors ($\tau=0$), the near stationary case and the close to local-to-unity process. The idiosynchratic terms $\mathbf{v}_{i,t}$ and $\epsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units. We start with the in the literature established case of stationary factors. In the small $T$ cases, $T=10$ and $T=20$, we first observe an overselection of the number of CAs, followed by an underselection. This pattern resembles the results in Margaritella and Westerlund, (2023) and is consistent across all values for $N$ and $\tau$ [^4] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$ If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors. Once the number of cross-sectional units and time periods passes 50, all four criteria select close to or exactly the correct number of factors.
@@
To investigate the behavior of the IC over different levels of non-stationarity, we increase $\tau$ in $0.05$ steps from 00 to $0.95$. Figure 1 presents the share of misselected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$. We also allow the idiosyncratics in $\mathbf{x}_{i,t}$ and $\mathbf{v}_{i,t}$ to be uncorrelated over time ($\rho=\rho_{v}=0$, in blue) and weakly correlated ($\rho=\rho_{v}=0.5$, in red). The share of misselected number of CAs remains relatively flat and below $20\%$ until $\tau$ reaches levels of around $0.5$, which is the midpoint. Neither of IC is strongly affected by weak dependence over time in the idiosyncratics. However, as shown in the Supplement, if the idiosynchratics in $\mathbf{X}_{i}$ are non-stationary as well, the IC equally fail to identify the correct number of CAs, with and without the adjustment of $T^{1+\tau}$. This again stresses the requirement for balanced stationarity in $\mathbf{F}$ and $\mathbf{V}_{i}$.
@@ 4 Conclusion
As predicted by the theory and our simulation results, the inability of IC to accommodate non-stationary factors is evident. The number of CAs is severely underselected, independent of the number of time periods and units. The underselection has severe consequences in terms of our examples. For the CCE estimator, since the number of factors is larger than the number of CAs, the rank condition fails and therefore consistency of the CCE estimator as well. In DCE case, the bootstrap procedure cannot help accommodate different factors in $\mathbf{X}$ and $\mathbf{y}$ anymore. In forecasting, incorrectly estimated factors harm the forecast comparison procedures.
@@ 5.3 Part $M\subset M_{0}$
Next, in case of $M\subset M_{0}$, we have that $\mathrm{rk}\left(\frac{1}{T^{1+\tau}}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)=m$ even asymptotically, and hence
@@ 5.4 Part $M_{0}\subset M$
with the obvious definitions of $\overline{\mathbf{H}}_{M_{0}}$ and $\overline{\mathbf{H}}_{M_{0}^{C}}$ and $\mathrm{rk}(\overline{\mathbf{H}}_{M})=g$. In what follows, the post-multiplication by this matrix leads to
@@
In what follows, we will use a slightly different path than De Vos and Stauskas, (2024) in order to sharpen some rates. Let us use the fact that $\mathbf{q}_{M_{0}}$ gives the minimal set of averages that asymptotically span the space of $\mathbf{F}_{\mathbf{x}}$. This means that $\mathrm{rk}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=\mathrm{rk}(\overline{\boldsymbol{\Gamma}}_{M_{0}})=m$ and this matrix is $m\times m$. This means that

### v2

@@ 2 Econometric Setup
This discussion suggests that accounting for non-stationarity of $\mathbf{F}$ in the objective function should solve the issue, but this is not the case. Assuming that $\tau$ is known, we normalize the objective function by $T^{1+\tau}$ as Assumption 1 (a) suggests. In the Supplement, we show that under $M\subset M_{0}$:
@@
$\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}=\ln\mathrm{det}\left[\mathbf{I}_{k}+\underbrace{\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x}}}}-\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}})\mathbf{X}_{i}}_{\mathbf{I}=O_{p}(1)}\left(\underbrace{\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}}\mathbf{V}_{i}}_{\mathbf{II}=O_{p}(T^{-\tau})}\right)^{-1}\right]+o_{p}(1),$ (2.7)
where the determinant does not converge to a strictly positive constant, again. The reason behind this is that the rank condition is satisfied in $\mathbf{II}$, and by following the steps of Karabiyik et al., (2017) and Westerlund, (2018), it admits (2.6). However, since idiosyncratics are stationary, “over-normalization” causes singularity, and we get a new imbalance between $\mathbf{I}$ and $\mathbf{II}$. Therefore, IC fundamentally depend not only on stationarity of factors, but also on stationarity of idiosyncratics. We confirm this with the additional simulations in the Supplement, where this adjustment becomes more harmful with $\tau$. If $M_{0}\subset M$, (2.5) and normalization by $T^{1+\tau}$ imply that (2.7) is still $\ln\det(\mathbf{I}_{k}+O_{p}(T^{-\tau}C_{N,T}^{-2})\times O_{p}(T^{\tau}))=O_{p}(C_{N,T}^{-2})$, which again means that $p_{N,T}$ needs to have different properties depending on whether $M_{0}\subset M$ or $M\subset M_{0}$.
@@ 3 Simulation Study
The data generating process of the simulation follows MW:
@@
where $\mathbf{X}_{i}=\left[\mathbf{x}_{i,1},...,\mathbf{x}_{i,T}\right]^{\prime}$ is a $T\times k$ matrix of observable regressors, $\boldsymbol{\beta}=\iota^{\prime}0.5$ a $k\times 1$ vector of unobserved parameters. The factor loadings are generated as $\boldsymbol{\Gamma}_{i}=[\mathbf{I}_{m},\mathbf{0}_{m\times(k-m)}]\psi_{i}$, that is, a matrix $m\times k$, with $\psi_{i}\sim N(1,1)$ and the elements of the vector $1\times m$ $\boldsymbol{\gamma}_{i}$ are drawn from $N(1,1)$. In the case of time- and space-independent errors, $\mathbf{e}_{i,t}$ is drawn from $N(\mathbf{0}_{k+1},\mathbf{I}_{k+1})$. In the case of errors correlated weakly across time and cross-sections we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
[…]
where $\boldsymbol{\epsilon}_{t}$ is a $N\times 1$ stack of $\epsilon_{i,t}\sim N(0,1)$ over $i$, and $\mathbf{w}_{i}$ is the $i$-th row ($1\times N$) of $\mathbf{W}$, a weight matrix with the $J$-th off diagonal elements equal to zero. $\rho$ is set to 0.5, $\kappa=0.2$ and $J=5$. The idiosyncratics in $\mathbf{X}_{i}$, $\mathbf{V}_{i}$, are also drawn as potentially weakly correlated over time and cross-sections. Here, $\mathbf{\Upsilon}_{t}$ is a $N\times k$ matrix, which stacks $\boldsymbol{\upsilon}_{i,t}\sim N(\mathbf{0}_{k},\mathbf{I}_{k})$ over $i$, and it is independent of $\boldsymbol{\epsilon}_{t}$. The vector $\boldsymbol{\iota}_{N,i}\in\mathbb{R}^{N\times 1}$ contains 1 in the $i$-th coordinate, and zeros elsewhere. $\rho_{v}=0$ for the time IID case and $\rho_{v}=0.5$ for weakly time-correlated idiosyncratics. In both cases, we allow for weak correlation across cross-sections, $\kappa_{v}=0.2$, $J_{v}=5$ and $\tilde{\mathbf{w}}_{i}$ is defined as $\mathbf{w}_{i}$ above. Therefore, we also test correlated idiosyncratics over time, whereas our theory prescribes uncorrelated innovations for $\tau\in(0,1)$. The factors are drawn as $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t}$, where
$\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{Q}T^{-\tau},\quad\mathbf{Q}=diag(q_{1},...,q_{m}),q_{j}\sim U(0,2)\quad\mathbf{u}_{f,t}\sim N\left(\mathbf{0}_{m},\sqrt{\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}}\right).$ (3.3)
@@
Table 6.6.5 below presents the average of the number of CAs selected for the case of stationary factors ($\tau=0$), the near stationary case and the close to local-to-unity process. The idiosynchratic terms $\mathbf{v}_{i,t}$ and $\epsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units. The Supplement presents additional results with weak correlation over time. We start with the in the literature established case of stationary factors. In the small $T$ cases, $T=10$ and $T=20$, we first observe an overselection of the number of CAs, followed by an underselection. This pattern resembles the results in MW and is consistent across all values for $N$ and $\tau$ [^4] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$ If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors. Once the number of cross-sectional units and time periods passes 50, all four criteria select close to or exactly the correct number of factors.
@@
Figure 1 presents the share of misselected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$, where we increase $\tau$ in $0.05$ steps from 00 to $0.95$. We also allow the idiosyncratics in $\mathbf{x}_{i,t}$ and $\mathbf{v}_{i,t}$ to be uncorrelated over time ($\rho=\rho_{v}=0$, in blue) and weakly correlated ($\rho=\rho_{v}=0.5$, in red). The share of misselected number of CAs remains relatively flat and below $20\%$ until $\tau$ reaches levels of around $0.5$, which is the midpoint. Neither of IC is strongly affected by weak dependence over time in the idiosyncratics. However, as shown in the Supplement, if the idiosynchratics in $\mathbf{X}_{i}$ are non-stationary as well, the IC equally fail to identify the correct number of CAs, with and without the adjustment of $T^{1+\tau}$. This again stresses the requirement for balanced stationarity in $\mathbf{F}$ and $\mathbf{V}_{i}$.
@@
4 Illustration: a Discussion of Serlenga and Shin, (2007)
MW apply $IC_{1}^{MW}$ and $IC_{2}^{MW}$ to revisit the study of Serlenga and Shin, (2007) (SS), who investigate the gravity equation of bilateral trade flows among 15 European countries ($N=91$ trading pairs) over 1960–2001 by the means of the CCE estimator. While SS include CAs arbitrarily, MW motivate their usage of IC as a theoretically backed and data-driven selection method. In neither of the studies are the time series properties of the variables examined.
Variables of interest include $\mathbf{z}_{i,t}=(\textit{Trade}_{i,t},\textit{RER}_{i,t},\textit{TGDP}_{i,t},\textit{RLF}_{i,t},\textit{SIM}_{i,t},\textit{EEC}_{i,t},\textit{EMU}_{i,t})^{\prime}$, where $\textit{Trade}_{i,t}$ (the dependent variable) denotes a total bilateral trade flow for a trading pair $i$, $\textit{RER}_{i,t}$ is the log real exchange rate between the pair currencies, $\textit{TGDP}_{i,t}$ represents the economic mass of the partners, $\textit{RLF}_{i,t}$ measures differences in the relative factor endowments. Moreover, $\textit{SIM}_{i,t}$ is the logarithm of a relative economic size, while dummies $\textit{EEC}_{i,t}$ and $\textit{EMU}_{i,t}$ represent European Economic Community and the European Monetary Union, respectively. Finally, $\textit{RERT}_{t}$ (log real exchange rate between the EU currencies and the dollar), a constant ($1$) and a linear trend ($t$) are treated as observed factors. Therefore, $\overline{M}$ constitutes 10 CAs to select from, and MW select $\widehat{M}=\{1,t,\textit{RERT}_{t},\overline{Trade}_{t},\overline{TDGP}_{t},\overline{RER}_{t},\overline{\textit{EMU}}_{t}\}$ with $g=7$ to control for $\mathbf{F}$. $\widehat{M}$ minimizes the remaining CS correlation in the CCE residuals (see Table 1 in MW).
We examine the series in SS further and detect strong evidence of unit roots in $\mathbf{F}$. Table 2 presents the results of the panel unit root test of Reese and Westerlund, (2016), which is the CAs equivalent of PANIC of Bai and Ng, (2004). Although we find strong evidence in favor of stationary idiosyncratics (at $1\%$ based on the PMSB statistic), we do not reject the null of unit root factors in virtually all series. [^5] The dummies $\textit{EEC}_{i,t}$ and $\textit{EMU}_{i,t}$ are omitted as they are unlikely to follow a typical factor structure. While MW acknowledge sensitivity of the results, our theory only strengthens the fact that the CCE-estimated coefficients in their Table 1 should be interpreted with some caution, and perhaps alternative combinations of CAs with $g\leq 10$ can provide a clearer picture.
Series $Trade$ $TDGP$ $RER$ $RLF$ $SIM$ $RERT$
Unit Roots in Factors
Errors
Table 2: PANICCA of Reese and Westerlund, (2016). - null of unit root not rejected; - the null rejected. For each series, the rest are used as regressors.
@@ 5 Conclusion
Our theory and simulation results show the inability of IC to accommodate non-stationary factors. The number of CAs is severely underselected, independent of the number of time periods and units. Underselection has severe consequences in terms of our examples. For CCE, the number of factors may be larger than the selected number of CAs, which makes the estimator inconsistent. In DCE case, the bootstrap procedure cannot help accommodate different factors in $\mathbf{X}$ and $\mathbf{y}$ anymore. In forecasting, incorrectly estimated factors may harm the forecast comparison procedures.
@@ 6.3 Part $M\subset M_{0}$
Next, in case of $M\subset M_{0}$, we have that $\mathrm{rank}\left(\frac{1}{T^{1+\tau}}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)=m$ even asymptotically, and hence
@@ 6.4 Part $M_{0}\subset M$
with the obvious definitions of $\overline{\mathbf{H}}_{M_{0}}$ and $\overline{\mathbf{H}}_{M_{0}^{C}}$ and $\mathrm{rank}(\overline{\mathbf{H}}_{M})=g$. In what follows, the post-multiplication by this matrix leads to
@@
In what follows, we will use a slightly different path than De Vos and Stauskas, (2024) in order to sharpen some rates. Let us use the fact that $\mathbf{q}_{M_{0}}$ gives the minimal set of averages that asymptotically span the space of $\mathbf{F}_{\mathbf{x}}$. This means that $\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=\mathrm{rank}(\overline{\boldsymbol{\Gamma}}_{M_{0}})=m$ and this matrix is $m\times m$. This means that